If the likelihood of binominal distribution for a tagged order form is successful probability

is 0.1, what is the probability that there are three tagged order forms in the sample of four?

Required

a. Determining Given n = 4 and p = 0.1 P(X = 3)


1
Expert's answer
2021-07-16T13:43:35-0400

n=4, p=0.1


probability that there are three tagged order forms in the sample of four

"P(X=3)=^4C_3p^3(1-p)^{4-3}"

"=4\times (0.1)^3.(0.9)
\\
 =3.6\times 0.001

\\

 =0.0036"